# Neuro-Symbolic Entity Alignment via Variational Inference

### v1

@@ 4.2 Inference with interpretability
To predict new alignments, there are two approaches: using the symbolic model or the neural model. The symbolic model infers alignment probabilities with the optimized weights $w$. Due to scalability concerns, symbolic methods generally adopt a lazy inference strategy that only preserves the confident pairs implied by the neighbor structure during inference. On the other hand, the neural model computes similarity scores for all entity pairs $(e,e^{\prime})\in{\mathcal{H}}$ using the learned parameters $\theta$, generating a ranked candidate list for each entity.
The evaluation of these models thus differs. Symbolic models are generally evaluated by precision, recall, and F1-score for their binary outputs, while neural models are assessed using hit@k and mean reciprocal ranks (MRR) for their ranked candidate lists. Following the practices in Qi et al. (2021) and Liu et al. (2022), we unify the evaluation metrics by treating the recall metric of symbolic models as equivalent to hit@1, facilitating comparison with neural models.
To enhance the interpretability of predictions, we adapt the optimized symbolic model into an explainer. For any given entity pair, the explainer generates a set of supporting rule path pairs that justify their alignment, each associated with a confidence score indicating its significance. The explainer operates in two modes: (1) hard-anchor mode, which generates supporting paths only from prealigned pairs, and (2) soft-anchor mode, which includes paths from both prealigned and inferred pairs, providing more informative interpretations.
By integrating a breadth-first search algorithm (detailed in Section A.2), the explainer efficiently generates high-quality interpretations. For truly aligned pairs, it typically produces high-confidence interpretations, while for non-aligned pairs, the interpretations may result in an empty set (indicating no supporting evidence) or have low confidence scores. See Figure 3 for a visualized comparison.
@@ 5.1 Experimental settings
In this section, we conduct experiments to answer the following questions. RQ1: Can NeuSymEA outperform existing neural, symbolic, and neuro-symbolic methods in terms of alignment performance? RQ2: Can symbolic and embedding-based methods complement each other in our framework? RQ3: How does the incorporated embedding-based model affect the alignment performance? RQ4: How does NeuSymEA interpret the inferred entity pairs, and how is the effectiveness of interpretations with respect to the rule length?
Datasets. We utilize the multilingual DBP15K dataset, which consists of three cross-lingual knowledge graph (KG) pairs: ja-en, fr-en, and zh-en. Note that the original full version (Sun et al., 2017) of this dataset contains a significant number of long-tail entities, presenting challenges for GCN-based models in terms of sparsity and large size. Therefore, many recent works (Wang et al., 2018; Mao et al., 2020; Liu et al., 2022) employ a condensed version, where long-tail entities and their connected triples are removed. For a comprehensive evaluation, we use both versions. Detailed statistics for these datasets are provided in Section B.1. Each dataset is divided into training, validation, and test sets following a 5-fold cross-validation scheme, with a ratio of 2:1:7.
Baselines and metrics. Baseline models include six neural models – GCNAlign (Wang et al., 2018), AlignE, BootEA (Sun et al., 2018), RREA (Mao et al., 2020), Dual-AMN (Mao et al., 2021), LightEA (Mao et al., 2022), one symbolic models – PARIS (Suchanek et al., 2012), and two neuro-symbolic models – PRASE (Qi et al., 2021), EMEA (Liu et al., 2022). We use Hit@1, Hit@10, and MRR as the evaluation metrics. For PARIS and PRASE that have binary outputs, we report their recall as Hit@1, following Liu et al. (2022).
Hyperparameters. NeuSymEA has two key hyperparameters: the number of EM iterations and the threshold $\delta$ for selecting positive pairs from the symbolic model. We tune these hyperparameters and select the best values based on the validation set. The search space for $\delta$ is $\{0.6,0.7,0.8,0.9,0.95,0.98,0.99\}$, while the number of iterations is searched from 1 to 9.
@@ 5.2.1 Comparison with baselines
Table 1: Entity alignment results on DBP15K dataset. The suffixes ”-D” and ”-L” indicate the use of Dual-AMN and LightEA as the neural models. The results of RREA and EMEA are omitted on the full dataset due to an OOM (Out of Memory) error.
@@
Group 1: Results on the full DBP15K dataset
@@
Group 2: Results on the condensed DBP15K dataset
@@
To answer RQ1, RQ2 and RQ3, we compare NeuSymEA with baseline models on two versions of the DBP15K dataset: the full version (group1) and the condensed version (group2). Results are presented in Table 1. The results for PRASE, and EMEA on the condensed DBP15K are adopted from the original EMEA paper. We employ both Dual-AMN and LightEA as the neural models in our framework, denoted as NeuSymEA-D and NeuSymEA-L, respectively. These results lead to several key observations:
First, NeuSymEA surpasses both symbolic and KG embedding-based models. This can be attributed to the fact that it combines the capacity of both sides: 1) the ability to precisely infer confident pairs with rules by leveraging the multi-hop relational structures; and 2) the ability to effectively model entity representations in a unified space to jointly optimize the alignment of two KGs. NeuSymEA seamlessly combines both symbolic reasoning and neural representations in a principled variational framework, leading to improved performance.
Second, NeuSymEA outperforms both neuro-symbolic models, PRASE and EMEA. This improvement can be largely attributed to the model objective design in our framework. While PRASE and EMEA treat the symbolic and neural models as separate components, NeuSymEA integrates them within a principled variational EM framework, unifying their objective as the log-likelihood of the observed variables. This approach allows for the joint optimization of both components, ensuring they work together to maximize the log-likelihood. By iteratively refining both components through the EM algorithm, NeuSymEA achieves a more coherent and convergent solution, leading to superior performance, as shown by the empirical results.
Finally, NeuSymEA demonstrates superior robustness across both versions of the DBP15K dataset. Comparisons between two groups of results offer an interesting insight: embedding-based models experience significant performance degradation when moving from the condensed version to the full version of DBP15K (e.g., MRR of Dual-AMN decreases from 0.815 to 0.717 on ja-en), while symbolic models, in contrast, show improvements. We attribute this to two key factors: (1) Embedding-based models rely on entity-level matching, which is sensitive to dataset size. As the dataset grows, the number of similar entity embeddings increases, leading to reduced accuracy. The full version of DBP15K contains significantly more entities than the condensed version, exacerbating this effect. (2) Symbolic models, on the other hand, perform path-level matching. Their effectiveness is constrained more by substructure heterogeneity and sparsity than dataset size. The full version of DBP15K includes more relational triples, which enhances the rule-mining process, ultimately making the symbolic models more effective in this scenario. NeuSymEA, by combining symbolic reasoning with KG embeddings, mitigates the shortcomings of both approaches, making it robust to changes in dataset scale and structure. This synergy allows NeuSymEA to consistently outperform baselines in both sparse and dense settings.
@@ 5.2.2 Evolution of rules and embeddings
We study how rules and embeddings evolve and interact with each other during the EM steps, with results shown in Figure 2. Results in the left subplot indicate that in each EM iteration, the number of rule-inferred pairs grows consistently with high precision, implying that the embedding model continuously improves the inference performance of rules. These precise pairs, in turn, enhance the performance of the neural model. As shown in the right subfigure, the MRR of the neural model converges within a few iterations.
@@ 5.2.3 Interpretations by the Explainer
Figure 3: (Left) Probability density of the top supporting rule’s confidence; (Middle) Number of supporting rules (thresholded by confidence score) relative to the maximum rule lengths under the soft anchor mode; (Right) Number of supporting rules relative to the maximum rule lengths under the hard anchor mode.
@@
We investigate the interpretations generated by the explainer on the fr-en dataset. The left subfigure of Figure 3 shows the probability density of confidence scores for supporting rules (with a maximum rule length of 3) associated with entity pairs. Positive pairs are derived from the test set, while negative pairs are created by replacing one entity in each test pair with another randomly sampled entity. The distinct confidence distributions indicate that positive pairs generally have more evidence for alignment, which aligns with intuition. However, the probability density distribution also reveals that some positive pairs do not have high confidence scores. Upon further examination, we found that many test pairs are isolated, i.e., they lack directly aligned neighbors. Despite this, NeuSymEA successfully generates supporting rules for isolated pairs by exploiting multihop dependencies. In Table 2, we provide several examples of supporting rules and their associated confidence scores for the queried entity pairs.
To examine the impact of rule length on the explainer’s effectiveness, the middle and right subfigures in Figure 3 show the number of supporting rules for test positive pairs as the maximum rule length increases. Compared to hard anchor mode, the explainer in soft anchor mode generates more high-quality supporting rules by leveraging inferred pairs as complementary anchor pairs, mitigating the sparsity issue. We also observe that increasing the maximum rule length leads to more high-quality rules; however, the number of high-confidence rules grows more slowly than lower-confidence rules. This can be attributed to our method for calculating confidence: the logical deduction-based approach computes a rule’s confidence as the product of the confidences of its decomposed length-one sub-rules (as in Equation 11). For example, a rule with two sub-rules, each with confidence $0.8$, results in an overall confidence of $0.8\times 0.8=0.64$. Considering this, the confidence score tends to decrease when the rule length increases, thus increasing the maximum length tends to discover supporting rules with lower scores.
Figure 4: Alignment performance with varying percentages of pairs as training data. For brevity, we present only the Hit@1 metric, with comprehensive results available in Section B.2.
@@ 5.2.4 Robustness in low resource scenario
Figure 5 demonstrates the model performance under low-resource settings. As the percentage of training data decreases, all models experience noticeable drops in Hit@1 performance. Despite this, NeuSymEA (both NeuSymEA-L and NeuSymEA-D) exhibits remarkable robustness across all datasets, consistently outperforming other models even with limited data. Notably, with only 1% of pairs used as training data, NeuSymEA-L achieves a Hit@1 score exceeding 0.7 on fr-en, rivaling or even surpassing the performance of some state-of-the-art models trained on 20% of the data. While models such as LightEA, Dual-AMN, and BootEA show improvement with additional training data, they still fall short of NeuSymEA. In contrast, traditional models like GCNAlign and AlignE demonstrate significant struggles in low-resource settings, underscoring their heavier dependence on larger training datasets.
@@ 6 Related work
Neuro-symbolic reasoning on knowledge graphs. Knowledge graphs (KGs) inherently contain logical rules within the relational patterns of their edges. Reasoning with these patterns is crucial for KG tasks such as KG completion (Chen et al., 2024a; Cheng et al., 2022; 2023; Zhang et al., 2024; Liu et al., 2024; Fang et al., 2023) and error detection (Zhang et al., 2022; 2023b). Neuro-symbolic methods aim to merge symbolic reasoning with neural representation learning, leveraging the precision and interpretability of symbolic approaches alongside the scalability and high recall of neural methods. In KG completion task, Guo et al. (2016) employ horn rules to regularize the learning of KG embeddings; Cheng et al. (2022) and Cheng et al. (2023) model the rule-based predictions as distributions conditioned on the input relational sequences, and parameterize these distributions using a recurrent neural network; Qu et al. (2020) and Cheng et al. (2022) enhances rule grounding by augmenting the fact set using a pre-trained KG-embedding model; Qu & Tang (2019), Zhang et al. (2020) and Chen et al. (2024a) models the joint probability of neural model and the symbolic model with a Markov random field. Despite extensive studies of neuro-symbolic reasoning in KG completion(DeLong et al., 2024; Cheng et al., 2024), they only consider single-KG structures, thus cannot be directly adopted to entity alignment which requires consideration of inter-KG structures.
Entity alignment. Recent models have sought to combine symbolic and neural approaches for entity alignment. For instance, Qi et al. (2021) enhances probabilistic reasoning by utilizing KG embeddings, employing a KG-embedding model to measure similarities during both updating and inference processes. Liu et al. (2022) implements self-bootstrapping with pseudo-labeling in a neural framework, using rules to choose confident pseudo-labels. However, it relies solely on unit-length rules, which restricts its effectiveness for long-tail entities. Furthermore, neither of these approaches offers interpretability for the inferred alignments.

### v2

@@ 4.2 Inference with interpretability
To predict new alignments, there are two approaches: using the symbolic model or the neural model. The symbolic model infers alignment probabilities with the optimized weights $w$. Due to scalability concerns, symbolic methods generally adopt a lazy inference strategy that only preserves the confidently inferred pairs during inference. On the other hand, the neural model computes similarity scores for all entity pairs $(e,e^{\prime})\in{\mathcal{H}}$ using the learned parameters $\theta$, generating a ranked candidate list for each entity.
The evaluation of these models thus differs. Symbolic models are generally evaluated by precision, recall, and F1-score for their binary outputs, while neural models are assessed using hit@k and mean reciprocal ranks (MRR) for their ranked candidate lists. Following the practices in [14] and [24], we unify the evaluation metrics by treating the recall metric of symbolic models as equivalent to hit@1, facilitating comparison with neural models.
To enhance the interpretability of predictions, we adapt the optimized symbolic model into an explainer. For any given entity pair, the explainer generates a set of supporting rule path pairs that justify their alignment, each associated with a confidence score indicating its significance. The explainer operates in two modes: ❶ hard-anchor mode, which generates supporting paths only from prealigned pairs, and ❷ soft-anchor mode, which includes paths from both prealigned and inferred pairs, providing more informative interpretations.
By integrating a breadth-first search algorithm (detailed in Section˜A.3), the explainer efficiently generates high-quality interpretations. For truly aligned pairs, it typically produces high-confidence interpretations, while for non-aligned pairs, the interpretations may result in an empty set (indicating no supporting evidence) or have low confidence scores. See Figure˜4 for a visualized comparison.
@@ 5.1 Experimental settings
Datasets. Main experiments use the DBP15K dataset, comprising three cross-lingual KG pairs: JA-EN, FR-EN, and ZH-EN. The original full version [29] of DBP15K resembles real-world KGs, posing challenges for GCN-based models due to sparsity and scale. Recent GCN-based models [19, 16, 17, 24] remove low-degree entities to get a smaller version with higher average degree. For thorough evaluation, we utilize both full and condensed versions. Dataset statistics are provided in Section˜B.1. For additional experiments on large KGs, we employ OpenEA [30] and DBP1M.
Two different dataset split strategies are used in the EA literature: ❶ a 3:7 train/test split, and ❷ a 5-fold cross-validation scheme with a 2:1:7 ratio for training, validation, and test sets, as used in OpenEA [30]. We adopt the latter for all algorithms to ensure fair comparison.
Baselines and metrics. Baseline models include seven neural models – GCNAlign [19], AlignE, BootEA [16], RREA [17], Dual-AMN [18], LightEA [20], PEEA [31], one symbolic models – PARIS [12], and two neuro-symbolic models – PRASE [14], EMEA [24]. We use Hit@1, Hit@10, and MRR as the evaluation metrics. For PARIS and PRASE that have binary outputs, we report their recall as Hit@1, following [24]. For RREA, Dual-AMN, and LightEA, which offer both basic and iterative versions, we adopt the iterative ones due to their generally superior performance.
Hyperparameters. NeuSymEA involves two main hyperparameters: the number of EM iterations and the symbolic model’s threshold $\delta$ for selecting positive pairs. We tune them on the validation set, searching $\delta$ in 0.6,0.7,0.8,0.9,0.95,0.98,0.99{0.6,0.7,0.8,0.9,0.95,0.98,0.99} and the number of iterations from 1 to 9. As shown in Section˜5.3, NeuSymEA is robust to $\delta$ and converges quickly.
@@ 5.2.1 Comparison with baselines
Table 1: Entity alignment results on DBP15K dataset. The suffixes "-D" and "-L" indicate the use of Dual-AMN and LightEA as the neural models. The results of RREA and EMEA are omitted on the full dataset due to an OOM (Out of Memory) error.
@@
Results on the full DBP15K dataset
@@
Results on the condensed DBP15K dataset
@@
Table˜1 compares NeuSymEA with baseline models on two versions of the DBP15K dataset: the full and the condensed version. Results for PRASE and EMEA on the condensed DBP15K are sourced from the original EMEA paper. The results yield three key observations:
First, NeuSymEA surpasses both symbolic and neural models. By integrating symbolic reasoning with neural representations, it: 1) captures multi-hop relational structures across KGs using rules; and 2) learns effective entity representations to compute pairwise similarities. Second, NeuSymEA outperforms other neuro-symbolic models. This improvement can be largely attributed to the model objective design in our framework. While PRASE and EMEA treat the symbolic and neural models as separate components, NeuSymEA unifies them under a joint probability objective. This enables joint optimization via Variational EM, yielding a more coherent and convergent solution with superior performance. Finally, NeuSymEA demonstrates robustness across both full and condensed datasets. Comparisons between two groups of results offer an interesting insight: neural models experience significant performance degradation in the full version of DBP15K (e.g., MRR of Dual-AMN decreases from 0.815 to 0.717 on JA-EN), while symbolic models, in contrast, show improvements. We attribute this to their different matching mechanisms: (1) neural models rely on entity-level matching, which is sensitive to dataset size. The full DBP15K includes more low-degree entities, increasing similar embeddings and reducing precision. (2) Symbolic models use path-level matching, which is less affected by dataset size but vulnerable to substructure heterogeneity. The full DBP15K’s additional connections via long-tail entities enhance rule-mining, boosting symbolic model performance. By integrating symbolic reasoning with KG embeddings, NeuSymEA overcomes these limitations, ensuring robustness to variations in dataset scale and structure.
@@ 5.2.2 Evolution of rules and embeddings
We study how rules and embeddings evolve and interact with each other during the EM steps, with results shown in Figure˜2. Results in the left subplot indicate that in each EM iteration, the number of rule-inferred pairs grows consistently with high precision, implying that the embedding model continuously improves the inference performance of rules. These precise pairs, in turn, enhance the performance of the neural model. As shown in the right subfigure, the MRR of the neural model converges within a few iterations.
5.2.3 Scalability on large datasets
Figure 3: Scalability analysis on large scale KGs. (Left) Hit@1 alignment performance on large KGs. (Right) Per-second processed entities of neural and symbolic components on different scales of KGs.
Figure˜3 illustrate the scalability of NeuSymEA. We evaluated its hit@1 accuracy and runtime efficiency across datasets of varying entity sizes, namely DBP15K, OpenEA100K, and DBP1M.
Performance scalability. In the left subfigure, we compare NeuSymEA with LightEA - the only strong baseline capable of processing large-scale knowledge graphs with million-scale entities. The results demonstrate NeuSymEA’s superior scalability in terms of performance.
Runtime Scalability. We separately evaluate the runtime performance of NeuSymEA’s neural and symbolic reasoning components. For the neural component, efficiency (entities processed per second) increases with dataset size. We attribute this to higher GPU utilization on larger datasets, which enhances computational efficiency. Conversely, the symbolic component’s efficiency initially increases but then slightly decreases. Upon investigation, we found that this phenomenon is related to the multiprocessing and batch-processing implementation. For smaller datasets (e.g., DBP15K), the overhead from process initialization and termination is significant. As dataset size grows to OpenEA100K, this overhead becomes negligible relative to inference runtime, leading to an efficiency improvement. When moving from OpenEA100K to DBP1M, the per-second processed entities slightly decline as the quadratic complexity of inference computation dominates.
@@ 5.2.4 Interpretations by the explainer
Figure 4: (Left) Probability density of the top supporting rule’s confidence; (Middle) Number of supporting rules relative to the maximum rule lengths under the soft anchor mode; (Right) Number of supporting rules relative to the maximum rule lengths under the hard anchor mode.
We investigate the interpretations generated by the explainer on the FR-EN dataset in Figure˜4. The left subfigure displays the probability density of confidence scores for supporting rules associated with entity pairs. Positive pairs come from the test set, while negative pairs are generated by randomly replacing one entity in each positive pair. The distinct confidence distributions show that positive pairs generally have stronger alignment evidence, as expected. Upon further examination, we found that many test pairs are isolated, i.e., they lack directly aligned neighbors. Despite this, NeuSymEA successfully generates supporting rules for isolated pairs by exploiting multi-hop dependencies.
The middle and right subfigures examine the impact of rule length on the explainer’s effectiveness, showing the number of supporting rules for positive pairs as the maximum rule length increases. In soft anchor mode, the explainer produces more high-quality supporting rules than in hard anchor mode by using inferred pairs as complementary anchors, mitigating substructure sparsity. While longer maximum rule lengths yield more high-quality supporting rules, however, the long rules have a lower confidence distribution compared to short rules. This can be attributed to our method for calculating confidence: the logical deduction-based approach computes a rule’s confidence as the product of the confidences of its decomposed unit-length sub rules (as in Equation 11). For instance, a rule comprising two unit-length sub-rules, each with a confidence of 0.8, has a combined confidence of $0.8\times 0.8=0.64$. Thus, confidence scores typically decrease as rule length increases.
@@ 5.2.5 Robustness in low resource scenario
Table 3: Comprehensive results with different ratios of training data, with best shown in bold.
Dataset Model 1% 5% 10% 20%
H@1 H@10 MRR H@1 H@10 MRR H@1 H@10 MRR H@1 H@10 MRR
JA-EN AlignE 0.007 0.034 0.016 0.080 0.268 0.143 0.244 0.588 0.356 0.433 0.783 0.552
BootEA 0.010 0.040 0.021 0.379 0.683 0.481 0.468 0.779 0.573 0.530 0.829 0.631
GCNAlign 0.029 0.128 0.063 0.127 0.368 0.206 0.209 0.515 0.310 0.331 0.662 0.443
PARIS 0.145 - - 0.340 - - 0.450 - - 0.565 - -
PRASE 0.163 - - 0.432 - - 0.508 - - 0.580 - -
Dual-AMN 0.239 0.519 0.334 0.509 0.795 0.611 0.652 0.887 0.738 0.750 0.927 0.815
RREA 0.253 0.486 0.332 0.558 0.830 0.653 0.672 0.903 0.756 0.789 0.956 0.853
LightEA 0.291 0.514 0.363 0.627 0.806 0.689 0.714 0.874 0.771 0.778 0.911 0.828
EMEA 0.411 - 0.488 0.630 - 0.710 0.688 - 0.764 0.736 - 0.807
PEEA 0.242 0.519 0.333 0.490 0.785 0.589 0.612 0.834 0.679 0.703 0.912 0.777
NeuSymEA-D 0.481 0.684 0.550 0.692 0.855 0.749 0.742 0.895 0.796 0.835 0.953 0.879
NeuSymEA-L 0.632 0.779 0.683 0.733 0.870 0.781 0.773 0.900 0.818 0.858 0.954 0.894
FR-EN AlignE 0.008 0.040 0.019 0.127 0.408 0.217 0.347 0.733 0.475 0.457 0.821 0.580
BootEA 0.009 0.041 0.020 0.418 0.746 0.529 0.490 0.809 0.598 0.579 0.872 0.681
GCNAlign 0.027 0.119 0.058 0.133 0.388 0.215 0.215 0.539 0.321 0.325 0.688 0.446
PARIS 0.195 - - 0.401 - - 0.479 - - 0.584 - -
PRASE 0.227 - - 0.514 - - 0.575 - - 0.633 - -
Dual-AMN 0.293 0.631 0.407 0.598 0.886 0.703 0.717 0.928 0.797 0.793 0.954 0.854
RREA 0.289 0.583 0.389 0.628 0.895 0.725 0.717 0.932 0.796 0.789 0.956 0.853
LightEA 0.430 0.663 0.509 0.723 0.885 0.781 0.779 0.914 0.828 0.827 0.943 0.870
EMEA 0.480 - 0.565 0.677 - 0.757 0.727 - 0.802 0.773 - 0.841
PEEA 0.285 0.588 0.385 0.552 0.812 0.642 0.665 0.875 0.738 0.748 0.937 0.815
NeuSymEA-D 0.642 0.833 0.709 0.768 0.916 0.820 0.811 0.939 0.856 0.835 0.953 0.879
NeuSymEA-L 0.737 0.874 0.785 0.806 0.921 0.848 0.827 0.937 0.867 0.858 0.954 0.894
ZH-EN AlignE 0.006 0.033 0.016 0.127 0.368 0.206 0.296 0.635 0.407 0.474 0.806 0.587
BootEA 0.006 0.029 0.014 0.396 0.689 0.495 0.498 0.782 0.594 0.575 0.847 0.668
GCNAlign 0.041 0.155 0.080 0.147 0.396 0.229 0.225 0.519 0.323 0.335 0.653 0.442
PARIS 0.059 - - 0.333 - - 0.429 - - 0.543 - -
PRASE 0.241 - - 0.461 - - 0.522 - - 0.593 - -
Dual-AMN 0.375 0.666 0.480 0.582 0.830 0.672 0.676 0.892 0.755 0.756 0.918 0.816
RREA 0.316 0.564 0.403 0.605 0.858 0.696 0.686 0.901 0.765 0.760 0.934 0.823
LightEA 0.507 0.673 0.565 0.670 0.819 0.723 0.727 0.860 0.775 0.770 0.894 0.816
EMEA 0.517 - 0.591 0.665 - 0.738 0.706 - 0.777 0.748 - 0.815
PEEA 0.288 0.586 0.388 0.532 0.801 0.622 0.649 0.871 0.710 0.725 0.905 0.790
NeuSymEA-D 0.589 0.750 0.645 0.704 0.856 0.757 0.763 0.897 0.809 0.815 0.926 0.855
NeuSymEA-L 0.676 0.799 0.720 0.735 0.858 0.779 0.773 0.882 0.811 0.804 0.934 0.841
Table˜3 demonstrates the model performance under low-resource settings. As the percentage of training data decreases, all models experience noticeable drops in Hit@1 performance. Despite this, NeuSymEA exhibits remarkable robustness across all datasets, consistently outperforming other models. Notably, with only 1% of pairs used as training data, NeuSymEA achieves a Hit@1 score exceeding 0.7 on FR-EN, rivaling or even surpassing the performance of some state-of-the-art models trained on 20% of the data.
5.3 Parameter analysis
We present the hit@1 performance of NeuSymEA across three datasets, varying hyperparameters, illustrated by a three-dimensional graph. The threshold hyperparameter $\delta$ is explored within the set {0.6, 0.7, 0.8, 0.9, 0.95, 0.98, 0.99}, while the number of EM iterations ranges from 1 to 9. Performance levels are indicated using a colormap. Performance sensitivity analysis in Figure˜5 reveals that for all datasets, performance generally improves as the iteration increases. On the other hand, the performance is less sensitive to the threshold $\delta$.
@@ 6 Related work
Neuro-symbolic reasoning on knowledge graphs. Neuro-symbolic methods aim to combine symbolic reasoning with neural representation learning, leveraging the precision and interpretability of symbolic approaches alongside the scalability and high recall of neural methods. In KG completion task, [32] and [33] employ horn rules to regularize the learning of KG embeddings; [34] and [28] model the rule-based predictions as distributions conditioned on the input relational sequences, and parameterize these distributions using a recurrent neural network; [26], [27] and [35] models the joint probability of the neural model and the symbolic model with a Markov random field, and employ gradient descent for weight updates. Despite extensive advancements of neuro-symbolic reasoning in KG completion, these studies only consider single-KG structures, thus cannot be directly adopted to entity alignment which requires consideration of inter-KG structures.
Entity alignment. Recent models have sought to combine symbolic and neural approaches for entity alignment. For instance, [14] enhances probabilistic reasoning with KG embeddings to measure entity-level and relation-level similarities. [24] implements self-bootstrapping with pseudo-labeling in a neural framework, using rules to choose confident pseudo-labels. However, it relies solely on unit-length rules, which restricts its effectiveness for long-tail entities. In contrast, we employ logic deduction to scale symbolic reasoning with long rules of any length. Unlike existing work that separately handles two reasoning components, our work models the unified joint probability of symbolic and neural inference within a markov random field. The symbolic component captures intra-KG and inter-KG structures using weighted logic rules, while the neural model learn expressive patterns in the embedding space.
7 Limitations
NeuSymEA is currently designed for EA between two KGs. Extending it to align multiple KGs simultaneously may require iterative pairwise alignments, which could be inefficient. A more sophisticated optimization paradigm is needed to adapt NeuSymEA for scalable multi-KG alignment.